OpenStudy (anonymous):

Okay wonderful Open Study peoples! If c(x)=4x^3-5x^2+2 and d(x)=3x^2+6x-10, find the value of c(y^3). Same as last time,medals up for grabs!!!

OpenStudy (rogue):

Just substitute y^3 in place of x for C(x).

OpenStudy (anonymous):

I know but......

OpenStudy (rogue):

\[c(x)=4x^3-5x^2+2\]\[c(y^3)=4(y^3)^3-5(y^3)^2+2\]

OpenStudy (rogue):

When you do a power to a power, just multiply the powers.\[c(y^3)=4(y^3)^3-5(y^3)^2+2 = c(y^3)=4y^9-5y^6+2\]